OpenStudy (s3a):

How is velocity an intensive property? My book says velocity is independent of mass so, why is the following wrong?: E = 1/2 * m * v^2 E = mc^2 So larger v means larger E and since c is constant, larger m.

OpenStudy (s3a):

Here is the problem.: Identify which of the following are extensive properties and which are intensive properties: (a) 10 m^3 volume, (b) 30 J of kinetic energy, (c) a pressure of 90 kPa, (d) a stress of 1000 kPa, (e) a mass of 75 kg, and (f) a velocity of 60 m/s. (g) Convert all extensive properties to intensive properties assuming m = 75 kg. Here is the solution.: (a) Extensive. If the mass is doubled, the volume increases. (b) Extensive. If the mass doubles, the kinetic energy increases. (c) Intensive. Pressure is independent of mass. (d) Intensive. Stress is independent of mass. (e) Extensive. If the mass doubles, the mass doubles. (f) Intensive. Velocity is independent of mass. (g) V/m = 10/75 = 0.1333 m^3 /kg E/m = 30/75 = 0.40 J/kg m/m = 75/75 = 1.0 kg/kg